# ATOM: AdapTive and OptiMized dynamic temporal knowledge graph construction using LLMs

### v1

@@ Abstract
In today’s rapidly expanding data landscape, knowledge extraction from unstructured text is vital for real-time analytics, temporal inference, and dynamic memory frameworks. However, traditional static knowledge graph (KG) construction often overlooks the dynamic and time-sensitive nature of real-world data, limiting adaptability to continuous changes. Moreover, recent zero- or few-shot approaches that avoid domain-specific fine-tuning or reliance on prebuilt ontologies often suffer from instability across multiple runs, as well as incomplete coverage of key facts. To address these challenges, we introduce ATOM [^1] The code, prompts, and dataset are available at https://github.com/AuvaLab/itext2kg. ATOM is available as an open-source Python library. (AdapTive and OptiMized), a few-shot and scalable approach that builds and continuously updates Temporal Knowledge Graphs (TKGs) from unstructured texts. ATOM splits input documents into minimal, self-contained “atomic” facts, improving extraction exhaustivity and stability. Then, it constructs atomic TKGs from these facts while employing a dual-time modeling that distinguishes when information is observed from when it is valid. The resulting atomic TKGs are subsequently merged in parallel. Empirical evaluations demonstrate that ATOM achieves $\sim 18\%$ higher exhaustivity, $\sim 17\%$ better stability, and over 90% latency reduction compared to baseline methods, demonstrating a strong scalability potential for dynamic TKG construction.
@@
Yassir LAIRGI1,2, Ludovic MONCLA1, Khalid BENABDESLEM1, Rémy CAZABET1,  and Pierre CLÉAU2 1LIRIS, INSA Lyon, Université Claude Bernard Lyon 1, France 2GAUC, Lyon, France {ludovic.moncla, khalid.benabdeslem, remy.cazabet}@liris.cnrs.fr {yassir.lairgi, pierre.cleau}@auvalie.com
@@ 1 Introduction
KG construction involves identifying entities, relationships, and attributes from diverse data sources to create structured knowledge representations. Traditionally, many approaches have focused on static Knowledge Graphs (KGs), which provide snapshots of knowledge without incorporating temporal dynamics. However, as real-world phenomena are inherently dynamic, static KGs, rarely or never updated, struggle to remain relevant and accurate (Jiang et al., 2023). In contrast, Temporal Knowledge Graphs (TKGs) integrate time dimensions by associating timestamps or time intervals with facts (e.g., (Einstein, was awarded, the Nobel Prize, in 1921)), making them particularly well-suited for analyzing changes, trends, and enabling temporal reasoning.
GraphRAG (Edge et al., 2024) and agent-based architectures (Xi et al., 2023) have demonstrated the potential of TKGs in retrieving and modeling dynamic information (Wu et al., 2024). Additionally, TKGs have been effectively used to model the memory of agents within agentic systems (Anokhin et al., 2024), highlighting their role in capturing the evolving nature of knowledge for adaptive and responsive systems.
Traditional methods for KG construction, often reliant on entity recognition and relation extraction, face several limitations. They typically depend on predefined ontologies and supervised learning techniques that require extensive human annotation (Al-Moslmi et al., 2020). Recent advances in Large Language Models (LLMs) (Jin et al., 2023) and zero- or few-shot techniques (Zhang et al., 2024; Carta et al., 2023; Hu et al., 2023) have paved the way for more flexible KG construction approaches that reduce dependency on extensive training datasets.
@@
In this paper, we propose ATOM (AdapTive and OptiMized), a few-shot and scalable dynamic TKG construction approach from unstructured text, ensuring stability and exhaustivity. ATOM introduces a strategy that decomposes unstructured text into atomic facts. Rather than processing these atomic facts sequentially, ATOM proposes an architecture with parallel 5-tuple extraction, followed by a parallel atomic merging mechanism. In the rest of the paper, we present related work in Section 2, our proposed approach ATOM in Section 3, experimental evaluation in Section 4, conclusion in Section 5, and present the limitations in Section 6.
@@ 2 Related work
Current zero- and few-shot approaches to KG construction, such as AttacKG+ (Zhang et al., 2024), iterative LLM prompting pipelines (Carta et al., 2023), LLM-Tikg (Hu et al., 2023), LLM Builder  [^2] https://llm-graph-builder.neo4jlabs.com/, and LLM Graph Transformer [^3] https://python.langchain.com/docs/how_to/graph_constructing/ aim to build KGs without requiring task-specific training. However, these methods suffer from inconsistencies such as unresolved entities and relations.
@@
AriGraph (Anokhin et al., 2024) integrated semantic and episodic memories to support reasoning, planning, and decision-making in LLM agents. However, their entity resolution method leads to semantic drift in the temporal KG, where, for example, a reference to “Apple” might ambiguously denote either the company or the fruit. Furthermore, scalability becomes problematic as the volume of unstructured data increases.
@@ 3.1 Problem statement
ATOM incorporates dual-time modeling, differentiating between when facts are observed and the temporal information conveyed by the facts themselves, which is characterized by a validity period. This approach better reflects real-world data (Rasmussen et al., 2025; Chekol and Stuckenschmidt, 2018; Meijer, 2022). This separation ensures TKG dynamism and proper inference of relative times by providing the observation time as context to the LLM (eg, ’month ago’).
@@
A fact in this snapshot is a 5-tuple (quintuple ) $(e_{s},r_{p},e_{o},t_{start},t_{end})$ indicating the relation $r_{p}\in\mathcal{R}^{t}$ holds between the subject entity $e_{s}\in\mathcal{E}^{t}$ and the object entity $e_{o}\in\mathcal{E}^{t}$. Technically, $t_{start}$ and $t_{end}$ are chosen to be lists to aggregate start and end validity timestamps to track the history of the same fact. Validity start and end timestamps can be unknown, in which case their respective lists are empty (denoted as $[.]$).
@@
Concretely, $\mathcal{G}^{t}_{i}$ is the set of 5-tuples $(e_{s},r_{p},e_{o},t_{start},t_{end})$ derived from a single atomic fact $f_{t,i}$ at observation time $t$.
@@
Figure F.1 in the Appendices illustrates a detailed example of ATOM’s pipeline.
@@ 3.2.1 Module-1: Atomic fact decomposition
ATOM does not construct TKGs directly from raw input documents but first decomposes them into atomic facts (Figure 1). This decomposition addresses a fundamental limitation of LLMs: the "forgetting effect" where models prioritize salient information in longer contexts while omitting key relationships, leading to incomplete knowledge extraction Liu et al. (2023). Following (Hosseini et al., 2024; Chen et al., 2023; Raina and Gales, 2024), ATOM uses LLM-based prompting for decomposition with an optimal chunk size to maintain high exhaustivity (determined experimentally in Section 4.3). However, while prior work focused on information retrieval applications, ATOM applies atomic decomposition specifically for TKG construction. This strategy addresses both conditions: it ensures exhaustivity (C1) by preventing information loss that occurs when LLMs process complex, multi-fact paragraphs, and enhances stability (C2) by providing clear, unambiguous contexts that reduce output variance across multiple runs. Each atomic fact is related to an observation time, and it is necessary to encapsulate the relative validity period presented in the context. The primary computational challenge of this approach is scale: a single document can yield hundreds or thousands of atomic facts. Sequential processing of each fact for 5-tuple extraction, followed by entity/relation and temporal resolution, becomes time-consuming. To address this challenge, ATOM employs a parallel architecture for both extraction and merging phases, as detailed in the subsequent modules.
@@ 2. RQ2:
    How does ATOM’s atomic fact decomposition ensure stability, exhaustivity, and improve the quality of the 5-tuples?
@@ 4.1 Metrics
To assess (C1), we adopt the metrics Hallucination (HALL), Omission (OM), and Match (MATCH) introduced by Ghanem and Cruz (2024). Given a gold-standard KG, MATCH denotes correctly extracted triplets, OM refers to triplets that exist in the source but are missing from the extracted KG, and HALL represents unsupported triplets. These definitions are extended from factual triplets $(e_{s},r_{p},e_{o})$ to temporal 5-tuples $(e_{s},r_{p},e_{o},t_{start},t_{end})$. A 5-tuple may exhibit correctness at the factual level while introducing temporal hallucination or omission. Hence, for 5-tuples whose factual components match the gold standard, the following definitions are established:
  • $\text{MATCH}_{t}$: 5-tuples whose $(t_{start},t_{end})$ is also correct.
  • $\text{HALL}_{t}$: 5-tuples containing temporal values not present in the gold standard.
  • $\text{OM}_{t}$: 5-tuples whose the $(t_{start},t_{end})$ are in the gold standard but are not reproduced in the extraction.
@@
The exhaustivity of the factual component is measured as:
$R_{\text{MATCH}}=\frac{|\text{MATCH}|}{|\text{MATCH}|+|\text{OM}|}$ (6)
The exhaustivity of the temporal component $R_{\text{MATCH}_{t}}$ is computed analogously with $|\text{MATCH}_{t}|$ in the numerator.
@@
It is assessed through the factual/temporal hallucination and omission rates:
$R_{\text{HALL}}=\frac{|\text{HALL}|}{|\text{MATCH}|+|\text{HALL}|}$ (7)
$R_{\text{OM}}=\frac{|\text{OM}|}{|\text{MATCH}|+|\text{OM}|}$ (8)
The $R_{\text{OM}_{t}}$ is computed similarly to $R_{\text{OM}}$ with $|\text{OM}_{t}|$ in the numerator. And $R_{\text{HALL}_{t}}$ is simply $R_{\text{MATCH}}$ - $R_{\text{MATCH}_{t}}$ - $R_{\text{OM}_{t}}$.
@@
It is measured using the cosine similarity between the centroid embeddings of 5-tuple sets obtained across independent runs. Let $\boldsymbol{c}^{(1)}$ denote the centroid vector obtained during the baseline run (RUN 1) and let $\boldsymbol{c}^{(r)}$ denote the centroid obtained at repetition $r$. Formally, the stability score is computed as:
$S_{r}=\cos\bigl(\boldsymbol{c}^{(1)},\,\boldsymbol{c}^{(r)}\bigr)=\frac{\bigl\langle\boldsymbol{c}^{(1)},\,\boldsymbol{c}^{(r)}\bigr\rangle}{\lVert\boldsymbol{c}^{(1)}\rVert\,\lVert\boldsymbol{c}^{(r)}\rVert}$ (9)
This score is computed for $r=2,3$.
@@
For entity/relation resolution, the false discovery rate (1-precision) is defined in (Lairgi et al., 2024). Since they overlook recall and F1-score, we extend the evaluation to include precision, recall, and F1-score for both entity resolution (ER) and relation resolution (RR), denoted as Metric-ER and Metric-RR, respectively. For temporal resolution, qualitative comparison is provided.
@@ 4.2 Datasets and baseline methods
DocRed (Yao et al., 2019) is unsuitable for temporal extraction due to inconsistencies (Tan et al., 2022), and TempDocRed (Zhu et al., 2025), though temporally enriched, focuses only on event start dates and named entities. CS-GS and Music-GS (Kabal et al., 2024), also used in iText2KG evaluations, lack temporal data. Therefore, the NYT News dynamic and temporal dataset (Singh, 2023), containing lead paragraphs from two million news articles since 2000, is adopted. From this, the 2020-COVID-NYT subset comprising 1,076 articles that focus on COVID-19 dynamics during 2020 is extracted. This subset is enriched with human-verified atomic facts and 5-tuples. Publication dates are used as observation dates (details on observation time modeling in Section D in the Appendices). More information about this dataset is provided in Table T.1 in the Appendices. To the best of our knowledge, an approach for resolving duplicate entities and relations while maintaining KG consistency among the SOTA methods for zero- and few-shot KG construction is supported only by iText2KG and Graphiti. The temporal aspect is handled by Graphiti only. Hence, it is the primary comparator of ATOM. In all experiments, the temperature is set to 0 to reduce hallucinations. text-embedding-large-3444https://platform.openai.com/docs/models/text-embedding-3-large is used for embeddings. $\theta_{E}=0.8$ and $\theta_{R}=0.7$ are estimated following Lairgi et al. (2024) (details are in Section C in the Appendices).
@@ 4.3 Exhaustivity deterioration in longer contexts
Exhaustivity is evaluated by iteratively concatenating lead paragraphs (increasing context size) and testing five SOTA LLMs claude-sonnet-4-2025-01-31555https://docs.claude.com/en/docs/about-claude/models/overview, gpt-4o-2024-11-20666https://platform.openai.com/docs/models/gpt-4o-2024-11-20, mistral-large-2411777https://docs.mistral.ai/getting-started/models/models_overview/, gpt-4.1-2025-04-14888https://openai.com/index/gpt-4-1/, o3-mini-2025-01-31999https://openai.com/index/openai-o3-mini/. Figure 2 shows a clear "forgetting effect". A decreased factual and temporal exhaustivity as token count increases across all models except claude-Sonnet-4-2025-01-31, which maintains the highest exhaustivity for atomic facts but degrades for 5-tuples. This indicates that LLMs prioritize salient facts in longer texts. Moreover, all models show higher exhaustivity for atomic fact decomposition than 5-tuple extraction. Atomic facts require surface-level parsing, while 5-tuples demand deeper semantic understanding of entities, relationship identification, and temporal extraction ($t_{start}$, $t_{end}$). This added complexity causes greater information loss. To mitigate information loss in the atomic fact decomposition, we empirically determine the optimal chunk size at <400<400 tokens to keep the exhaustivity >0.8>0.8. For subsequent evaluations, we use claude-sonnet-4-2025-01-31 for atomic fact decomposition based on its superior performance. For 5-tuple extraction, both gpt-4.1-2025-04-14 and claude-sonnet-4-2025-01-31 perform comparably, hence gpt-4.1-2025-04-14 is used due to its lower cost. Section 4.5 evaluates exhaustivity gains from using atomic facts as input.
Figure 2: Exhaustivity vs. token count as context for (a) the atomic fact decomposition (b) 5-tuples extraction.

### v2

@@ Abstract
In today’s rapidly expanding data landscape, knowledge extraction from unstructured text is vital for real-time analytics, temporal inference, and dynamic memory frameworks. However, traditional static knowledge graph (KG) construction often overlooks the dynamic and time-sensitive nature of real-world data, limiting adaptability to continuous changes. Moreover, recent zero- or few-shot approaches that avoid domain-specific fine-tuning or reliance on prebuilt ontologies often suffer from instability across multiple runs, as well as incomplete coverage of key facts. To address these challenges, we introduce ATOM [^1] The code, prompts, and dataset are available at https://github.com/AuvaLab/itext2kg. ATOM is available as an open-source Python library. (AdapTive and OptiMized), a few-shot and scalable approach that builds and continuously updates Temporal Knowledge Graphs (TKGs) from unstructured texts. ATOM splits input documents into minimal, self-contained “atomic” facts, improving extraction exhaustivity and stability. Then, it constructs atomic TKGs from these facts, employing a dual-time modeling that distinguishes between when information is observed and when it is valid. The resulting atomic TKGs are subsequently merged in parallel. Empirical evaluations demonstrate that ATOM achieves $\sim 18\%$ higher exhaustivity, $\sim 33\%$ better stability, and over 90% latency reduction compared to baseline methods, demonstrating a strong scalability potential for dynamic TKG construction.
@@
Yassir LAIRGI1,2, Ludovic MONCLA1, Khalid BENABDESLEM1, Rémy CAZABET1,  and Pierre CLÉAU2 1INSA Lyon, CNRS, UCBL, LIRIS, UMR5205, 69621 Villeurbanne, France 2GAUC, Lyon, France {ludovic.moncla, khalid.benabdeslem, remy.cazabet}@liris.cnrs.fr {yassir.lairgi, pierre.cleau}@auvalie.com
@@ 1 Introduction
KG construction involves identifying entities, relationships, and attributes from diverse data sources to create structured knowledge representations. Traditionally, many approaches have focused on static KGs, which provide snapshots of knowledge without incorporating temporal dynamics. However, as real-world phenomena are inherently dynamic, static KGs, rarely or never updated, struggle to remain relevant and accurate (Jiang et al., 2023). In contrast, Temporal Knowledge Graphs (TKGs) integrate time dimensions by associating timestamps or time intervals with facts (e.g., (Einstein, was awarded, the Nobel Prize, in 1921)), making them particularly well-suited for analyzing changes, trends, and enabling temporal reasoning.
GraphRAG (Edge et al., 2024) and agent-based architectures (Xi et al., 2025) have demonstrated the potential of TKGs in retrieving and modeling dynamic information (Wu et al., 2024). Additionally, TKGs have been effectively used to model the memory of agents within agentic systems (Anokhin et al., 2025), highlighting their role in capturing the evolving nature of knowledge for adaptive and responsive systems.
Despite this potential, deploying Dynamic TKGs in production imposes latency constraints. As new information arrives continuously, systems must update the graph in real-time. In such contexts, sequential approaches often become prohibitive bottlenecks, creating a critical limitation for scalable, real-time analytics (Bian, 2025).
Traditional methods for KG construction, often reliant on entity recognition and relation extraction, face several limitations. They typically depend on predefined ontologies and supervised learning techniques that require extensive human annotation (Al-Moslmi et al., 2020). Recent advances in Large Language Models (LLMs) (Jin et al., 2024) and zero- or few-shot techniques (Zhang et al., 2024; Carta et al., 2023; Hu et al., 2024) have paved the way for more flexible KG construction approaches that reduce dependency on extensive training datasets.
@@
In this paper, we propose ATOM (AdapTive and OptiMized), a few-shot and scalable dynamic TKG construction approach from unstructured text, ensuring stability and exhaustivity. ATOM introduces a strategy that decomposes unstructured text into atomic facts. Rather than processing these atomic facts sequentially, ATOM proposes an architecture with parallel 5-tuple extraction, followed by a parallel atomic merging mechanism. In the remainder of the paper, we present related work in Section 2, our proposed approach, ATOM, in Section 3, experimental evaluation in Section 4, a conclusion in Section 5, and limitations in Section 6.
@@ 2 Related work
Current zero- and few-shot approaches to KG construction, such as AttacKG+ (Zhang et al., 2024), iterative LLM prompting pipelines (Carta et al., 2023), LLM-Tikg (Hu et al., 2024), LLM Builder  [^2] https://llm-graph-builder.neo4jlabs.com/, and LLM Graph Transformer [^3] https://python.langchain.com/docs/how_to/graph_constructing/ aim to build KGs without requiring task-specific training. However, these methods suffer from inconsistencies such as unresolved entities and relations.
@@
AriGraph (Anokhin et al., 2025) integrated semantic and episodic memories to support reasoning, planning, and decision-making in LLM agents. However, their entity resolution method leads to semantic drift in the temporal KG, where, for example, a reference to “Apple” might ambiguously denote either the company or the fruit. Furthermore, scalability becomes problematic as the volume of unstructured data increases.
@@ 3.1 Problem statement
ATOM incorporates dual-time modeling, differentiating between when facts are observed and the temporal information conveyed by the facts themselves, which is characterized by a validity period. This approach better reflects real-world data (Rasmussen et al., 2025; Chekol and Stuckenschmidt, 2018; Meijer, 2022). This separation enhances TKG dynamism and proper inference of relative times by providing the observation time as context to the LLM (eg, ’month ago’).
@@
A fact in this snapshot is a 5-tuple (quintuple ) $(e_{s},r_{p},e_{o},t_{\text{start}},t_{\text{end}})$ indicating the relation $r_{p}\in\mathcal{R}^{t}$ holds between the subject entity $e_{s}\in\mathcal{E}^{t}$ and the object entity $e_{o}\in\mathcal{E}^{t}$. Technically, $t_{\text{start}}$ and $t_{\text{end}}$ are chosen to be lists to aggregate start and end validity timestamps to track the history of the same fact. Validity start and end timestamps can be unknown, in which case their respective lists are empty (denoted as $[.]$).
@@
Concretely, $\mathcal{G}^{t}_{i}$ is the set of 5-tuples $(e_{s},r_{p},e_{o},t_{\text{start}},t_{\text{end}})$ derived from a single atomic fact $f_{t,i}$ at observation time $t$.
@@
Figure 2 illustrates a detailed example of ATOM’s pipeline.
Figure 2: Example overview of ATOM’s pipeline. It begins with atomic fact decomposition, followed by the extraction of atomic TKGs from these facts, which are then merged in parallel. When an incoming update arrives, ATOM handles the temporal resolution by transforming the end action into the affirmative part while modifying only the $t_{\text{end}}$, then merges the resulting atomic TKG with the existing DTKG.
@@ 3.2.1 Module-1: Atomic fact decomposition
ATOM does not construct TKGs directly from raw input documents but first decomposes them into atomic facts (Figure 1). This decomposition addresses a fundamental limitation of LLMs: the "forgetting effect" where models prioritize salient information in longer contexts while omitting key relationships, leading to incomplete knowledge extraction Liu et al. (2024). Following (Hosseini et al., 2024; Chen et al., 2024; Raina and Gales, 2024), ATOM uses LLM-based prompting for decomposition with an optimal chunk size to maintain high exhaustivity (determined experimentally in Section 4.3). However, while prior work focused on information retrieval applications, ATOM applies atomic decomposition specifically for TKG construction. This strategy addresses both conditions: it enhances exhaustivity (C1) by preventing information loss that occurs when LLMs process complex, multi-fact paragraphs, and enhances stability (C2) by providing clear, unambiguous contexts that reduce output variance across multiple runs. Each atomic fact is related to an observation time, and it is necessary to encapsulate the relative validity period presented in the context. The primary computational challenge of this approach is scale: a single document can yield hundreds or thousands of atomic facts. Sequential processing of each fact for 5-tuple extraction, followed by entity/relation and temporal resolution, becomes time-consuming. To address this challenge, ATOM employs a parallel architecture for both extraction and merging phases, as detailed in the subsequent modules.
@@ 2. RQ2:
    How does ATOM’s atomic fact decomposition enhance stability, exhaustivity, and improve the quality of the 5-tuples?
@@ 4.1 Metrics
To assess (C1), the performance is evaluated at two levels: factual and factual-temporal. Given a gold-standard KG, factual true positives ($\text{TP}_{f}$) denote correctly extracted triplets, false negatives ($\text{FN}_{f}$) refer to triplets that exist in the source but are missing from the extracted KG (omission), and false positives ($\text{FP}_{f}$) represent unsupported triplets (hallucination). These definitions are extended to the factual-temporal level for 5-tuples $(e_{s},r_{p},e_{o},t_{\text{start}},t_{\text{end}})$. A 5-tuple is considered valid only if it exhibits correctness at both the factual and temporal levels. Hence, the factual-temporal true positives $\text{TP}_{f,t}$ are 5-tuples whose $t_{\text{start}}$ and $t_{\text{end}}$ are both correct. Consequently, $\text{TP}_{f,t}\subseteq\text{TP}_{f}$.
@@
The exhaustivity or recall is reported at two levels of strictness. The exhaustivity of the factual component:
$R_{f}=\frac{|\text{TP}_{f}|}{|\text{TP}_{f}|+|\text{FN}_{f}|}$ (6)
The exhaustivity of the factual-temporal component $R_{f,t}$ is computed analogously with $|\text{TP}_{f,t}|$ in the numerator. It measures the proportion of the total gold standard 5-tuples that are both factually and temporally correct.
@@
It is assessed through Recall, Precision, and $F_{1}$-score. The precision of the factual component:
$P_{f}=\frac{|\text{TP}_{f}|}{|\text{TP}_{f}|+|\text{FP}_{f}|}$ (7)
The precision of the factual-temporal component $P_{f,t}$ is computed similarly to $P_{f}$ with $|\text{TP}_{f,t}|$ in the numerator. It measures the proportion of all extracted 5-tuples that are both factually and temporally correct. The $F_{1,f}$ and $F_{1,f,t}$ scores are computed as the harmonic means of their respective precision and recall metrics.
@@
It is measured using the Jaccard similarity between the 5-tuple sets obtained across independent runs. Let $\boldsymbol{s}^{(1)}$ denote the set of 5-tuples obtained during the baseline run (RUN 1) and let $\boldsymbol{s}^{(r)}$ denote the set of 5-tuples obtained at repetition $r$. Formally, the stability score $S_{r}$ is computed as:
$S_{r}=J\bigl(\boldsymbol{s}^{(1)},\,\boldsymbol{s}^{(r)}\bigr)=\frac{\bigl\lvert\boldsymbol{s}^{(1)}\cap\boldsymbol{s}^{(r)}\bigr\rvert}{\bigl\lvert\boldsymbol{s}^{(1)}\cup\boldsymbol{s}^{(r)}\bigr\rvert}$ (8)
This score is computed for repetitions $r=2,3$.
@@
For entity/relation resolution, the false discovery rate (1-precision) is defined in (Lairgi et al., 2024). Since they overlook recall and $F_{1}$-score, we extend the evaluation to include precision ($P$), recall ($R$), and $F_{1}$-score for both entity resolution (ER) and relation resolution (RR), denoted as $\text{Metric}_{ER}$ and $\text{Metric}_{RR}$, respectively. For temporal resolution, a qualitative comparison is provided.
@@ 4.2 Datasets and baseline methods
Identifying a suitable and publicly available dataset for evaluating temporal extraction remains a significant challenge due to the task’s specific structural requirements. While DocRed (Yao et al., 2019) is a standard benchmark for relation extraction, it is largely unsuitable for temporal analysis owing to documented labeling inconsistencies (Tan et al., 2022). Although TempDocRed (Zhu et al., 2025) introduces temporal layers, its scope remains limited, focusing exclusively on named entities and event start dates. Furthermore, the CS-GS and Music-GS datasets (Kabal et al., 2024), lack the necessary temporal dimensions. Consequently, this study adopts the NYT News dynamic and temporal dataset (Singh, 2023), which provides extensive temporal coverage through two million lead paragraphs spanning from 2000 to the present. News articles provide diverse temporal dynamics, including events with specific start and end times, evolving situations, and relative temporal expressions that are representative of many real-world scenarios requiring TKG construction. From this dataset, the 2020-COVID-NYT subset comprising 1,076 articles that focus on COVID-19 dynamics during 2020 is extracted. This subset is enriched with human-verified atomic facts and 5-tuples. Publication dates are used as observation dates. Comprehensive details regarding the annotation guidelines, observation time modeling, and dataset statistics are provided in the Appendices (Sections E, D, and Table T.1). To the best of our knowledge, an approach for resolving duplicate entities and relations while maintaining KG consistency among the SOTA methods for zero- and few-shot KG construction is supported only by iText2KG and Graphiti. The temporal aspect is handled by Graphiti only. Hence, it is the primary comparator of ATOM. In all experiments, the temperature is set to 0 to prefer deterministic outputs and minimize stochasticity. text-embedding-large-3444https://platform.openai.com/docs/models/text-embedding-3-large is used for embeddings. $\theta_{E}=0.8$ and $\theta_{R}=0.7$ are estimated following Lairgi et al. (2024) (details are in Section C in the Appendices).
@@ 4.3 Exhaustivity deterioration in longer contexts
Exhaustivity is evaluated by iteratively concatenating lead paragraphs (increasing context size) and testing five SOTA LLMs claude-sonnet-4-2025-01-31555https://docs.claude.com/en/docs/about-claude/models/overview, gpt-4o-2024-11-20666https://platform.openai.com/docs/models/gpt-4o-2024-11-20, mistral-large-2411777https://docs.mistral.ai/getting-started/models/models_overview/, gpt-4.1-2025-04-14888https://openai.com/index/gpt-4-1/, o3-mini-2025-01-31999https://openai.com/index/openai-o3-mini/. Figure 3 shows a clear "forgetting effect". A decreased factual and factual-temporal exhaustivity as token count increases across all models except claude-Sonnet-4-2025-01-31, which maintains the highest exhaustivity for atomic facts but degrades for 5-tuples. This indicates that LLMs prioritize salient facts in longer texts. Moreover, all models show higher exhaustivity for atomic fact decomposition than 5-tuple extraction. Atomic facts require surface-level parsing, while 5-tuples demand deeper semantic understanding of entities, relationship identification, and temporal extraction ($t_{\text{start}}$, $t_{\text{end}}$). This added complexity causes greater information loss. To mitigate information loss in the atomic fact decomposition, we empirically determine the optimal chunk size at <400<400 tokens to keep the exhaustivity >0.8>0.8. For subsequent evaluations, we use claude-sonnet-4-2025-01-31 for atomic fact decomposition based on its superior performance. For 5-tuple extraction, both gpt-4.1-2025-04-14 and claude-sonnet-4-2025-01-31 perform comparably, hence gpt-4.1-2025-04-14 is used due to its lower cost. Section 4.5 evaluates exhaustivity gains from using atomic facts as input.
Figure 3: Exhaustivity vs. token count as context for (a) the atomic fact decomposition (b) the 5-tuple extraction. The x-axis displays the context size (lead paragraphs) only; total input size includes additional system prompt tokens ($\sim$1.4k for (a) and $\sim$1.7k for (b)).